4) A right triangle is a triangle with one angle measuring 90°. In a right triangle, the sides are related by Pythagorean Theorem, where c is the hypotenuse (the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 3 feet and 4 feet. Please
Found 2 solutions by checkley71, marshalltamika:
Answer by checkley71(8403)

a~2+b~2=c~2 or 3~2+4~2=c~2 or 9+16=c~2 25=c~2 or c-sqrt 25 or c=5

Answer by marshalltamika(3)

Since you are given two sides of the triangle and you know that neither of the two are the hypotenuse.
You can say that either a=3 and b=4 or the other way where a=4 and b=3. Then by using the pythagorean theorem which states a^2+b^2=c^2 you will be able to get your answer.

So
(3)^2+(4)^2=c^2
9+16=c^2
25=c^2
Now by taking the square root of each side you get that c=+ - 5, but since you are dealing with lengths and you can't have a negative length, then c=5.